OpenStudy (anonymous):

The driveway to a house has the shape of a square with a side of 8 m adjacent to a rectangle with width 4.5 m and length 14 m. How much will it cost to pave the driveway, if the cost per square foot to pave is $22.5?

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

\(\sf\Large $22.5 \times ((8\times 8) +(4.5\times 14))\)

OpenStudy (anonymous):

1503

TheSmartOne (thesmartone):

hmm how did you get that? Make sure to do the stuff in the parenthesis first before.

TheSmartOne (thesmartone):

before doing the rest*

OpenStudy (anonymous):

28,57.5

TheSmartOne (thesmartone):

Correct.

OpenStudy (anonymous):

okay, my answer book states $31,750 is the correct answer

TheSmartOne (thesmartone):

Tricky question, aha. They gave us the dimensions in meter, and the price in square feet.

OpenStudy (anonymous):

ikr!!! lol, ive been working on this one fore awhile

OpenStudy (anonymous):

lets see if @robtobey can help?

TheSmartOne (thesmartone):

So basically, (8 x8) + (4.5 x 14) = 63+ 64 = 127 meters squared using this conversion to make it feet squared... and then so 1367.02 x 22.5 = 30757.95

TheSmartOne (thesmartone):

That gives us a closer answer, but not the same exact answer .-.

OpenStudy (anonymous):

@robtobey can ya help?

OpenStudy (anonymous):

1 square meter = 10.7639104 square ft. http://www.bing.com/search?q=convert%20square%20meters%20to%20square%20feet&pc=cosp&ptag=A316A5D3C6E&form=CONBDF&conlogo=CT3210127 127 x 10.7639 x 22.5 = 30757.84425

OpenStudy (anonymous):

thankyou!! guys!!

OpenStudy (anonymous):

Your welcome.